40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 485     Entry time: Sun May 18 18:44:48 2008
Author: rana 
Type: Summary 
Category: SUS 
Subject: Optical Lever SUM Trend - 80 days 
I used the OL-Trend.xml dataviewer template to make this plot. Looks like the ETMY optical lever
slowly degraded over the last few months and then finally died 10 days ago. Would someone please
replace this laser and tune the lens position to minimize the spot size on the quad?
Attachment 1: e.pdf  10.024 MB  | Hide | Hide all
e.pdf
ELOG V3.1.3-